# Wrong Signs are Alright

### v1

@@ I Introduction
In this work, we show that loop effects due to low-energy modes that travel on closed future-directed trajectories cause divergent backreaction on the classical configurations. The reason for these divergences is that the loop modes are on-shell. Even before the closed future-directed trajectory appears, slightly off-shell closed trajectories contribute to loops, causing increasingly strong backreaction. These are new, configuration-dependent divergences and they cannot be removed by standard, background-independent counterterms.
@@ III Vacuum energy
Subtracting this divergence is impossible if one wishes to make the theory finite around the vacuum or any causal configurations. Indeed, the localized backgrounds we are considering can be described as coherent states excited from the vacuum. For this reason, the available counterterms are fixed by renormalizing $n$-point vacuum-to-vacuum correlators, which are not affected by the divergence discussed above.
Therefore, this divergence indicates that for a system coupled to gravity or with generic self-interactions, the low-energy description breaks down and the configuration becomes UV sensitive. This invalidates the assumption that the classical time-machine configuration can be described at low energy.
@@ IV Strong Coupling
We start studying quantum corrections to the two-point correlator for $\varphi$ evaluated at points that are away from any closed trajectories. Focusing for simplicity on the quartic self-interaction, at one loop we have:
@@
where $G_{i,x}\equiv G_{\varphi}(x_{i},x)$ and the derivatives are contracted with a dimensionless, constant tensor that we omit. In this expression, the integral over space-time will always lead to divergent contributions in which the loop propagator is evaluated on closed on-shell trajectories: $G_{\varphi}(x,x)\sim 1/\tilde{g}^{\mu\nu}k_{\mu}k_{\nu}$. Other large contributions will come from slightly off-shell closed trajectories that are formed before the closed on-shell trajectories appear. These divergences depend on where the closed trajectories are located with respect to $x_{1},x_{2}$, and involve both IR and UV modes. As in the case of the previous section, these effects cannot be subtracted by renormalization. Indeed the counterterms needed would make the theory infinite in any configuration with a standard causal structure. Moreover, note that these effects would be present even if one had canonically normalized $\varphi$, since the background is weakly coupled and there is no appreciable screening. At the level of power counting the 4-volume of the closed curves scales as $L^{2}\sigma_{\gamma}$, meaning that these contributions are enhanced by $c_{2}/\sigma_{\gamma}^{2}$.
@@ V Time Machines Decay
This condition is easily satisfied e.g. if $k=q$. Momentum conservation in this process will impose $k^{\prime}-q^{\prime}=(q-k)_{x\to y}$, this last being the parallel transported momentum $k-q$ along the internal line labeled by $p$ from one interaction point to the other. The corresponding amplitude can be obtained by applying LSZ to the two-point correlator:
@@ VI Discussion
In this work, we have shown how quantum effects protect the causal structure of quantum field theory at low energy, regardless of faster-than-light propagation. Backreaction from low-energy quanta on closed future-directed trajectories diverges, probing in turn the UV dynamics and making these configurations unpredictive. The same effect leads to very large backreaction from slightly off-shell quanta before a closed on-shell trajectory is formed, making the preparation of time-machine configurations impossible in the regime of validity of the low-energy theory. These divergences cannot be subtracted if the closed trajectories are created starting from a configuration with the usual causal structure of special relativity.
We described three different divergent loop effects: a divergent contribution to the energy-momentum tensor, the strong coupling of excitations propagating around time machine backgrounds and the decay of the classical configuration in radiation. Although we framed our discussion in terms of a specific theory for a shift-symmetric scalar, our analysis applies as long as the closed future-directed trajectories can be described in the geometric optics approximation. Moreover, since the divergences depend uniquely on having on-shell loop contributions from closed future-directed trajectories, the same will hold beyond our present approximation.

### v2

@@ I Introduction
In this work, we show that loop effects due to low-energy modes that travel on closed future-directed trajectories cause divergent backreaction on the classical configurations. The reason for these divergences is that the loop modes are on-shell, meaning that their presence is independent on the regularization technique employed. Even before the closed future-directed trajectory appears, slightly off-shell closed trajectories contribute to loops, causing increasingly strong backreaction. These are new, kinematic and configuration-dependent divergences and they cannot be removed by standard counterterms while keeping the theory predictive.
@@ III Vacuum energy
Subtracting this divergence is impossible if one wishes to make the theory predictive and finite around the vacuum or any causal configurations. Infact, our result in Eq. (9) corresponds to a kinematic-dependent divergence in $n$-point correlators. This is because the localized classical backgrounds we are considering can be described as coherent states excited from the vacuum. This means that the expectation value on the classical configuration can be rewritten as a sum of vacuum expectation values of multiple field operators. The momentum space kinematics of these field operators is determined by the background, e.g. by the size and relative speed of the blobs. Therefore, the divergence in Eq. (9) originates as a kinematic dependent divergence in the vevs that resum to the expectation value on the classical configuration. In particular, for kinematics that describe slow/distant enough blobs with no classical time-machine trajectory, all the vevs will be finite once the counterterms of the theory are fixed. Instead, changing the kinematics of the fields, one will encounter the closed on-shell trajectories and the divergence exposed above. This situation is in close analogy with the singularity encountered in GR as a result of gravitational collapse with initial conditions (kinematics) such that a black hole is formed. The same theory with initial conditions that differ only by the kinematics of the field excitations, will have no black hole formation and no singularity, making it unphysical to remove the singularity in the black hole interior by adding ad-hoc, configuration dependent counterterms. Another similar case is what happens in gravity-mediated scattering in 4 dimensions. There, the forward limit scattering of e.g. two particles is plagued by divergences that do not appear at finite momentum-exchange, and that cannot be consistently removed by counterterms. In the case of time-machine configurations, the divergence encountered in Eq. (9) indicates that for a system coupled to gravity or with generic self-interactions, the low-energy description breaks down and the configuration becomes UV sensitive. This invalidates the assumption that the classical time-machine configuration can be described consistently at low energy.
@@
In this section, our use of the point-splitting regulator has only been a tool to derive a simple relation between stress-energy tensor and two-point function. The results discussed here do not depend on the regularization technique that one would use to compute explicitly Eq. (9), as the divergent loop contribution can be regularized but not subtracted.
@@ IV Strong Coupling
We start studying quantum corrections to the two-point correlator for $\varphi$ evaluated at points that are away from any closed trajectories. In particular, we restrict our analysis to space-time points that can be time-ordered with respect to each other and with respect to the compact time-machine configuration. For instance, one could choose points that are in time-slices that do not intersect the time-machine configuration. This restriction allows to compute correlators e.g. in the in-out formalism. Indeed, despite the presence of the time-machine, one can define time-ordering through the chosen foliation of space-time and compute in-out correlators using the $\varphi$ Hamiltonian, which can be obtained considering the boosted blobs. Despite the separation between the chosen points and the closed trajectory, the latter still brings divergent virtual contributions to correlators, leading to strong coupling. Focusing for simplicity on the quartic self-interaction, at one loop we have:
@@
where $G_{i,x}\equiv G_{\varphi}(x_{i},x)$ with same $\epsilon$ prescription as Eq. (8) - and the derivatives are contracted with a dimensionless, constant tensor that we omit. In this expression, the integral over space-time will always lead to divergent contributions in which the loop propagator is evaluated on closed on-shell trajectories: $G_{\varphi}(x,x)\sim 1/\tilde{g}^{\mu\nu}k_{\mu}k_{\nu}$. Other large contributions will come from slightly off-shell closed trajectories that are formed before the closed on-shell trajectories appear. These divergences depend on where the closed trajectories are located with respect to $x_{1},x_{2}$, and involve both IR and UV modes. As in the case of the previous section, these effects cannot be subtracted by renormalization. Indeed the counterterms needed would make the theory infinite in any configuration with a standard causal structure. Moreover, note that these effects would be present even if one had canonically normalized $\varphi$, since the background is weakly coupled and there is no appreciable screening. At the level of power counting the 4-volume of the closed curves scales as $L^{2}\sigma_{\gamma}$, meaning that these contributions are enhanced by $c_{2}/\sigma_{\gamma}^{2}$.
@@
In alternative to our approach, one could study the system by computing in-in correlators and restricting to time-slices and time intervals that do not contain the time-machine. This strategy should still probe the effects we have discussed so far.
@@ V Time Machines Decay
This condition is easily satisfied e.g. if $k=q$. Momentum conservation in this process will impose $k^{\prime}-q^{\prime}=(q-k)_{x\to y}$, this last being the parallel transported momentum $k-q$ along the internal line labeled by $p$ from one interaction point to the other. The two on-shell modes will travel to null infinity along light-like geodesics of the acoustic metric that are not trapped in the locus of closed light-like geodesics. Similarly to before, considering such modes makes it possible to define a precise time ordering and to compute in-out correlators. Even more, we can apply LSZ to the two-point correlator to derive the matrix element for the production of two on-shell modes from the classical background configuration:
@@ VI Discussion
In this work, we have shown how quantum effects protect the causal structure of quantum field theory at low energy, regardless of faster-than-light propagation. Backreaction from low-energy quanta on closed future-directed trajectories diverges, probing in turn the UV dynamics and making these configurations unpredictive. The same effect leads to very large backreaction from slightly off-shell quanta before a closed on-shell trajectory is formed, making the preparation of time-machine configurations impossible in the regime of validity of the low-energy theory. These divergences are kinematic dependent and cannot be subtracted while keeping the theory finite and predictive, much like singularities arising in gravitational collapse or gravitational scattering.
We described three different divergent loop effects: a divergent contribution to the energy-momentum tensor, the strong coupling of excitations propagating away from the time machine backgrounds and the decay of the classical configuration in radiation. Although we framed our discussion in terms of a specific theory for a shift-symmetric scalar, our analysis applies as long as the closed future-directed trajectories can be described in the geometric optics approximation. Moreover, since the divergences depend uniquely on having on-shell loop contributions from closed future-directed trajectories, the same will hold beyond our present approximation.